Hi, Mine is a word problem.
The force of gravity on Jupiter is about 2 1/2 times that of Earth. How much would a 140 lb object weigh on Jupiter?
I thought the equation would look like this:
5/2x=140
But I dont remember what to do after

Answer by stanbon(75887)

The force of gravity on Jupiter is about 2 1/2 times that of Earth. How much would a 140 lb object weigh on Jupiter?
I thought the equation would look like this:
(5/2)x=140
-----
Jupiter weight = (2 1/2)*140 = (5/2)*140 = 350 lbs.
=======================
Cheers,
Stan H.
==============